OpenStudy (anonymous):

solve: 2sin²θ-sinθ=0 please show how you got your answer Thank you so much!!

OpenStudy (phi):

if the problem were 2 x^2 -x=0 could you find the values for x ?

OpenStudy (anonymous):

i got down to: 2cosθ=1 and cosθ=1/2

OpenStudy (phi):

you mean sin , don't you ? but can you solve 2 x^2 -x=0 ?

OpenStudy (anonymous):

yeah sorry

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

oh i forgot to add in the equation "-1" after the sin

OpenStudy (phi):

what do you get ?

OpenStudy (phi):

I would factor out an x from both terms as a start.

OpenStudy (anonymous):

okay so it would be 2x-1, right?

OpenStudy (phi):

or are you saying we have a different problem than what you posted ?

OpenStudy (anonymous):

and then x=0 and x=1/2

OpenStudy (anonymous):

the problem is: 2sin²θ-sinθ-1=0

OpenStudy (phi):

ok, so we want to solve 2x^2 -x -1 =0 If it factors it will be (2x 1 )(x 1 )=0 and we need to put in the correct signs the -1 at the end means the signs will be different the -x means the bigger one is minus can you figure it out ?

OpenStudy (anonymous):

so it will be (x-1) and (x+.5) which then will be x=1 and x=-.5

OpenStudy (anonymous):

and then i got cosθ=1 and cosθ=-.5 is this right, i tried using u-substitution because my teacher told me to

OpenStudy (phi):

if the problem is 2sin²θ-sinθ-1=0 then x is really sin of some angle so you have sin θ = 1 or sin θ = -0.5 I am not sure why you think it is cosine ??

OpenStudy (anonymous):

sorry i just got confused for a sec, but i know its sin

OpenStudy (phi):

one way to find the values of sin is plot it from 0 to 2pi (or 0 to 360 degrees)